OpenStudy (anonymous):

Find the distance between the points (-3, 11) and (5, 5).

jimthompson5910 (jim_thompson5910):

Label the two points (x1,y1) and (x2,y2)

jimthompson5910 (jim_thompson5910):

So x1 = -3, y1 = 11 and x2 = 5, y2 = 5

jimthompson5910 (jim_thompson5910):

Now evaluate x2 - x1 and tell me what you get

jimthompson5910 (jim_thompson5910):

x2 - x1 5 - (-3) ... replace x1 with -3 and x2 with 5 5 + 3 8 See how I got that?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

So (x2-x1)^2 = (8)^2 = 64

jimthompson5910 (jim_thompson5910):

Now what about y2 - y1 ???

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

what is y2 ?

OpenStudy (anonymous):

2\[\sqrt{10}\]

OpenStudy (anonymous):

y2?

jimthompson5910 (jim_thompson5910):

The two given points are (-3, 11) and (5, 5) In general, they are (x1,y1) and (x2,y2)

jimthompson5910 (jim_thompson5910):

See how y1 matches up with 11 ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what does y2 match with?

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

so y2 - y1 5 - 11 -6

jimthompson5910 (jim_thompson5910):

Then square it to get (-6)^2 = 36

jimthompson5910 (jim_thompson5910):

The next step is to take these squares, 64 and 36, and add them to get 64+36 = 100 Finally, take the square root sqrt(100) = 10

jimthompson5910 (jim_thompson5910):

So the distance between the two points is 10 units

OpenStudy (anonymous):

thankyou

jimthompson5910 (jim_thompson5910):

You can prove this visually by plotting the two points. Then drawing a circle centered at one point and that passes through the other point. You'll see that the radius of this circle is 10 units.

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

ok thanks